### v1

# Home Spaces and Semiflows for the Analysis of Parameterized Petri Nets

Abstract
After rapidly recalling basic notations of Petri Nets, home spaces, and semiflows, we focus on $\mathcal{F^{+}}$, the set of semiflows with non-negative coordinates where the notions of minimality of semiflows and minimality of supports are particularly critical to develop an effective analysis of invariants and behavioral properties of Petri Nets such as boundedness or even liveness. We recall known decomposition theorems considering semirings such as $\mathbb{N}$ or $\mathbb{Q^{+}}$, and then fields such as $\mathbb{Q}$. The decomposition over $\mathbb{N}$ is being improved with a necessary and sufficient condition.
Then, we regroup a number of properties (old and new) around the notions of home spaces and home states which in combination with semiflows are used to efficiently support the analysis of behavioral properties.
We introduce a new result on the decidability of liveness under the existence of a home state. We end this section with new results about the structure and behavioral properties of Petri Nets, illustrating again the importance of considering semiflows with non-negative coordinates.
As examples, we present two related Petri Net modeling arithmetic operations (one of which is an Euclidean division), illustrating how semiflows and home spaces can be used in analyzing the liveness of the parameterized model and underlining the efficiency brought by using minimal semiflows of minimal supports as well as the new results on the structure of the model.
Keywords:
Invariants Home spaces Home states Petri Nets Generating Sets Semiflows Boundedness Liveness.
1 Introduction
1.1 Motivations
Parallel programs, distributed digital systems, telecommunication networks, or cyber-physical systems are entities that are complex to design, model, and verify. Using formal verification at different stages of the system development life cycle is a strong motivation and provides us with the rationale for introducing the notions of semiflows and home spaces that are at the core of this paper. In this regard, invariants are of paramount importance as they are almost systematically used in system specifications to describe specific behavioral properties. One can argue that properties such as liveness, deadlock freeness, or boundedness are in some way invariants since they must hold regardless of the evolution of the digital system under study.
How can invariants be meaningfully combined to prove complex behavioral properties that one invariant alone cannot represent? In this regard, we will study a companion notion to invariants, namely home spaces, characterized as a set of always accessible markings. Often, behavioral properties can be proven starting from a given home space, especially when it is possible to use characteristics common to any marking of this home space.
Most of the time, engineers and researchers will be trying to prove that a formula that belongs to a system specification is an invariant, meaning that the formula holds during any possible evolution of their model. But, can we find a way by which invariants or at least a meaningful subset of invariants can be organized and concisely described, while some of them can be discovered by computation? Such invariants that do not belong in the system specification, can just express a sub-property of a more complex known one; however, they also can reveal an under-specified model or an unsuspected function of the system under study (which in turn, could constitute a component of a security breach). How can invariants be decomposed into simpler and verifiable properties? How to determine whether a given decomposition is more effective than another one with regard to formal verification?
In this paper, we provide some elements to answer these questions and show how basic arithmetic, linear algebra, or algebraic geometry can efficiently support invariant calculus. In such a setting, linear algebra can also be applied and utilized to prove a wide array of behavioral properties.
One of our motivations was to go beyond regrouping a number of known algebraic results dispersed throughout the Petri Nets literature, and introduce new results allowing to accurately position these results by considering semirings such as $\mathbb{N}$ or $\mathbb{Q^{+}}$, then fields such as $\mathbb{Q}$, especially regarding generating sets of semiflows ([Mem23]). Beyond proving that a formula is an invariant, can we find a way in which they can be organized or concisely described, a way in which they cas be discovered or computed? How can invariants be combined to represent meaningful behaviors? How can invariants be decomposed into simpler and verifiable properties? How to determine whether a given decomposition is more effective than another one with regard to formal verification? We will combine invariants and home spaces to address these questions, and illustrate how to proceed through two examples.
This paper can be considered as a continuation of the work started in [Mem23], providing new results particularly on home spaces as well as new examples. We want to show how linear algebra or algebraic geometry can efficiently sustain invariant calculus and can be applied and utilized to prove a large variety of behavioral properties, sometimes with simple arithmetic reasoning. When Petri Net are parameterized, this type of reasoning can be useful to determine in which domain of these parameters, behavioral properties can be satisfied.
1.2 Outline and contributions
After providing some basic notations in Section 2 and recalling a first set of classic properties for semiflows (in $\mathbb{Z}$ or $\mathbb{N}$) in Section 3, the notions of generating sets and minimality are briefly recalled from [Mem23] in Section 4.
The three decomposition theorems of Section 4.2 have been first published in [Mem78] then improved in [Mem23]. Here, the first theorem is extended once more to fully characterize minimal semiflows and generating sets over $\mathbb{N}$. The other two theorems are just recalled for completeness.
Then, the notion of home space is described Section 5 with a set of old and new results linked to their structure and later to their key relation with liveness in Section 5.2. In particular, a new decidability result is provided for Petri Nets with home states linked to Karp and Miller’s coverability tree finite construction.
Subsequently, Theorem 5.2 is new and describes three extremums regarding any semiflow and place in a support of a semiflow. This result can be computed from any generating set. These important details were never stressed out before despite their importance from a computational point of view, and their impact in analysis automation.
These results will be used in the analysis of two examples presented in Section 6 where two parameterized examples are given to illustrate how invariants and home spaces can be associated with basic arithmetic reasoning to prove behavioral properties of a Petri Net.
Section 7 concludes and provides a possible avenue for future research.
2 Basic notations
In this section, we briefly recall Petri Nets, including the notion of potential state space that is usual in Transition Systems, introducing notations that will be used in this paper. Then, we define semiflows in $\mathbb{Z}$ and basic properties in $\mathbb{N}$ highlighting why semiflows in $\mathbb{N}$ may be considered more useful to analyze behavioral properties.
[…]
A marking (or state in Transition Systems) $q:\ P\rightarrow\mathbb{N}$ allows representing the evolution of the system along the execution (or firing) of a transition $t$ or of a sequence of transitions $\sigma$ (i.e., a word in $T^{*}$). We say that $t$ is enabled at marking $q$ if and only if $q\geq Pre(\cdot,t),$ and as an enabled transition at $q$ (we sometimes write that $q\in\texttt{Dom}(t)$), $t$ can be executed, reaching a marking $q^{\prime}$ from $q$ such that:
$q^{\prime}=q+Pre(\cdot,t)+Post(\cdot,t).$
[…]
We also define $Q$, the set of all potential markings (also known as state space in Transition Systems). Without additional information on the domain in which marking of places may vary, we assume $Q=\mathbb{N}^{d}$
$RS(PN,Init)$ denotes the reachability set of a Petri Net PN from a subset $Init$ of $Q$: $RS(PN,Init)=\{q\in Q\ |\ \exists\ a\in Init,\ a\overset{*}{\rightarrow}q\}$. $RG(PN,Init)$, $LRG(PN,Init)$, and $LCT(PN,Init)$ will denote the corresponding reachability graph without labels for $RG(PN,Init)$ (as in Figure 1), with labels in $T$ for $LRG(PN,Init)$, while $LCT(PN,Init)$ is the covering tree.
3 Petri Nets and Semiflow basic properties
Definition 1 (Semiflow)
A Semiflow $f$ is a solution of the following homogeneous system of $|T|$ diophantine equations:
$f^{\top}Post(\cdot,t)=f^{\top}Pre(\cdot,t),\ \ \forall t\in T,$ (1)
where $x^{\top}y$ denotes the scalar product of the two vectors $x$ and $y$, since $f,Pre(\cdot,t)$ and $Post(\cdot,t)$ can be considered as vectors once the places of $P$ have been ordered.
$\mathcal{F}$ and $\mathcal{F}^{+}$ denote the sets of solutions of the system of equations (1) that have their coefficients in $\mathbb{Z}$ and in $\mathbb{N}$, respectively.
Considering a Petri Net $PN$ with its initial marking $q_{0}$ and the set of reachable markings from $q_{0}$ through all sequences of transitions denoted by $RS(PN,q_{0})$, any non-null solution $f$ of the homogeneous system of equations (1) allows to directly deduce the following invariant of $PN$ defined by its $Pre$ and $Post$ functions (used in the system of equations (1) that $f$ satisfies):
$\forall\ q\ \in RS(PN,q_{0}):f^{\top}q=f^{\top}q_{0}.$ (2)
In the rest of the paper, we abusively use the same symbol ‘0’ to denote $(0,...,0)^{\top}$ of $\mathbb{N}{{}^{n}}$, for all $n$ in $\mathbb{N}$. The support of a semiflow $f$ is denoted by $\left\|f\right\|$ and is defined by
$\left\|f\right\|=\{x\in P\ |\ f(x)\neq 0\}.$
We will use the usual component-wise partial order in which $(x_{1},x_{2},\dots,x_{d})^{\top}\leq(y_{1},y_{2},\dots,y_{d})^{\top}$ if and only if $x_{i}\leq y_{i}$, for all $i\in\{1,\dots,d\}$.
The most interesting set of semiflows, from a behavioral-analysis standpoint, is $\mathcal{F}^{+}$, defined over natural numbers. This can be seen through the three following properties. First, we define the positive and negative supports of a semiflow $f\in\mathcal{F}$ as:
$\left\|f\right\|_{+}=\left\{p\in P\ |\ f(p)>0\right\}$
and
$\left\|f\right\|_{-}=\left\{p\in P\ |\ f(p)<0\right\},$
with $\left\|f\right\|=\left\|f\right\|_{-}\cup\ \left\|f\right\|_{+}$. We can then rewrite Equation (2) as:
$f^{\top}q=\left|\sum_{p\in\left\|f\right\|_{+}}f(p)q(p)\right|-\left|\sum_{p\in\left\|f\right\|_{-}}f(p)q(p)\right|=f^{\top}q_{0}.$ (3)
As we can see, the formulation of Equation (3) is a subtraction between the weighted number of tokens in the places of the positive support and the weighted number of tokens in the places of the negative support of $f$. This expression allows deducing an invariant since, by Equations (3), it remains constant during the evolution of the Petri Net. A first general property can be immediately deduced by recalling that any marking $q$ belongs to $\mathbb{N}^{d}$ and that a subset $A$ of places is bounded .
Property 1
For any semiflow $f\in\mathcal{F}$, $\left\|f\right\|_{+}$ is bounded if and only if $\in\left\|f\right\|_{-}$ is bounded.
Of course, if $\left\|f\right\|_{-}=\varnothing$, then $f\in\mathcal{F}^{+}$ and $\left\|f\right\|$ is necessarily structurally bounded (i.e., bounded from any initial marking). More generally, considering a weighting function $f$ over $P$ being defined over non-negative integers and verifying the following system of inequalities:
$f^{\top}Post(\cdot,t)\leq f^{\top}Pre(\cdot,t),\ \ \forall t\in T,$ (4)
the following properties can be easily proven [Mem78]:
Property 2
If $f\geq 0$ is such that it verifies Equation 4, then the set of places of $\left\|f\right\|$ is structurally bounded.
Moreover, the marking of any place $p$ of $\left\|f\right\|$ has an upper bound:
$q(p)\leq\frac{f^{T}q_{0}}{f(p)},\ \ \forall q\in RS(PN,q_{0}).$
If $f>0$, then $\left\|f\right\|_{+}=\left\|f\right\|=P$, and the Petri Net is also structurally bounded. The reverse is also true: if the Petri Net is structurally bounded, then there exists a strictly positive solution for the system of inequalities above (see [Sif78] or [Bra82]). This property is actually false for a semiflow that satisfies Equation 1 but would have at least one negative element, and constitutes a first reason for particularly considering weight functions $f$ over $P$ being defined over non-negative integers including $\mathcal{F}^{+}$.
The following corollary can be directly deduced from the fact that any semiflow in $\mathcal{F}^{+}$ satisfies Property 2:
Corollary 1
For any place $p$ belonging to at least one support of a semiflow of $\mathcal{F}^{+}$, an upper bound $\mu$ can be defined for the marking of $p$ relatively to an initial marking $q_{0}$ such that:
$\forall q\in RS(PN,q_{0}),\ \ q(p)\leq\mu(p,q_{0})=\min_{\{f\in\mathcal{F}^{+}\ |\ f(p)\neq 0\}}\frac{f^{\top}q_{0}}{f(p)}.$
We will see with Theorem 5.2 that this bound is computable.
A second reason for particularly considering a semiflow $f$ as being defined over non-negative integers is that the system of inequalities
$f^{T}q_{0}\geq f^{T}Pre(\cdot,t),\ \ \ \ \forall t\in T,$ (5)
becomes a necessary condition for any transition $t$ to stand a chance to be enabled from any reachable marking from $q_{0}$, then to be live. In [Bra82], $f^{T}Pre(\cdot,t)$ is called the enabling threshold of $t$.
Property 3
If $t$ is a transition and $f\in\mathcal{F}^{+}\setminus\{0\}$ such that $f^{T}q_{0}<f^{T}Pre(\cdot,t)$, then $t$ cannot be executed from $\langle PN,q_{0}\rangle$.
This property can be interesting when the model is defined with parameters, since some values of these parameters for which the model is not live (see example of Figure 2) can be rapidly pruned away.
At last, the following known property ([Mem78], [Bra82], or [STC98]) can easily be proven true in $\mathcal{F}^{+}$ and not true in $\mathcal{F}$:
Property 4
If $f$ and $g$ are two semiflows with non-negative coefficients, then we have: $\left\|f+g\right\|=\left\|f\right\|\cup\left\|g\right\|$.
If $\alpha$ is a non-null integer then $\left\|\alpha f\right\|=\left\|f\right\|$.
This property is used to prove theorem 4.3 section 4.2 and theorem 5.2 section 5.3. These results have been cited and utilized many times in various applications going beyond computer science, electrical engineering, or software engineering. For instance, they have been used in domains such as population protocols [CEL23] or biomolecular chemistry relative to chemical reaction networks [JACB18], which brings us back to the C. A. Petri’s original vision, when he highlighted that his nets could be used in chemistry. Many other applications can be found in the literature.
4 Generating sets and minimality
The notion of generating sets for semiflows is well known and efficiently supports the handling of an important class of invariants. Several results have been published, starting from the initial definition and structure of semiflows [Mem77] to a wide array of applications used especially to analyze Petri Nets [CTSH03, DL16, JACB18, Wol19].
Minimality of semiflows and minimality of their supports are critical to understand how to best decompose semiflows. Invariants directly deduced from minimal semiflows relate to smaller weighted quantities of resources, simplifying the analysis of behavioral properties. Furthermore, the smaller the support of semiflows, the more local their footprint (i.e., the more constrained the potential exchanges between resources is). In the end, these two notions of minimality will foster analysis optimization.
4.1 Three definitions
Definition 2 (Generating set)
A subset $\mathcal{G}$ of $\mathcal{F^{+}}$ is a generating set over a set $\mathbb{S}$ (where $\mathbb{S}\in\{\mathbb{N},\mathbb{Q^{+}},\mathbb{Q}\}$ with $\mathbb{Q^{+}}$ denoting the set of non-negative rational numbers) if and only if for all $f\in\mathcal{F}^{+}$, we have $f=\sum_{g_{i}\in\mathcal{G}}\alpha_{i}g_{i}$, where $\alpha_{i}\in\mathbb{S}$ and $g_{i}\in\mathcal{G}$.
Since $\mathbb{N}\subset\mathbb{Q^{+}}\subset\mathbb{Q}$, a generating set over $\mathbb{N}$ is also a generating set over $\mathbb{Q^{+}}$, and a generating set over $\mathbb{Q^{+}}$ is also a generating set over $\mathbb{Q}$. However, the reverse is not true and is, in our opinion, a source of some inaccuracies that can be found in the literature (see [GV03], for instance). Therefore, it is important to specify over which set of {$\{\mathbb{N},\mathbb{Q^{+}},\mathbb{Q}\}$ the coordinates (used for the decomposition of a semiflow) vary.
Several definitions around the concept of minimal semiflow were introduced in [STC98], p. 319, in [CST03], p. 68, [KJ87], [CMPAW09], or in [Mem78, Mem83]. However, we will only consider two basic notions in order theory: minimality of support with respect to set inclusion and minimality of semiflow with respect to the component-wise partial order on $\mathbb{N}{{}^{d}}$, since the various definitions found in the literature as well as the results of this paper can be described in terms of these two classic notions.
Definition 3 (Minimal support)
A nonempty support $\left\|f\right\|$ of a semiflow $f$ is minimal with respect to set inclusion if and only if $\nexists\ g\in\mathcal{F}^{+}\setminus\{0\}$ such that $\left\|g\right\|\subset\left\|f\right\|$.
Definition 4 (Minimal semiflow)
A non-null semiflow $f$ is minimal with respect to $\leq$ if and only if $\nexists\ g\in\mathcal{F^{+}}\setminus\{0,f\}$ such that $g\leq f$.
A minimal semiflow cannot be decomposed as the sum of another semiflow and a non-null non-negative vector. This remark yields an initial insight into the foundational role of minimality in the decomposition of semiflows. We are looking for characterizing generating sets such that they allow analyzing various behavioral properties as efficiently as possible. That is to say that we want generating sets as small as possible and, at the same time, able to easily handle semiflows in $\mathcal{F}$. First, the number of minimal semiflows over $\mathbb{N}$ can be quite large. Second, considering a basis over $\mathbb{Q}$ is of course relevant to handle $\mathcal{F}$, while less when it is about $\mathcal{F}^{+}$, and may not capture behavioral constraints as easily. We will have to consider $\mathbb{Q}^{+}$.
4.2 Three decomposition theorems
Generating sets can be characterized thanks to three decomposition theorems. A first version of them can be found in [Mem78] with their proofs. A second version can be found in [Mem23] with improvements. Here, Theorem 4.1, which is valid over $\mathbb{N}$, is extended to a necessary and sufficient condition that characterizes a minimal semiflow and generating sets over $\mathbb{N}$. This result is provided with a new proof using Gordan’s lemma (see Lemma 1). Theorems 4.2 and 4.3 are recalled for completeness and are unchanged from [Mem23].
4.2.1 Decomposition over non-negative integers
The fact that there exists a finite generating set over $\mathbb{N}$ is non-trivial and is often taken for granted in the literature on semiflows. In fact, this result was proven by Gordan, circa 1885, then Dickson, circa 1913. Here, we directly rewrite Gordan’s lemma [AB86] by adapting it to our notations.
Lemma 1
(Gordan circa 1885) Let $\mathcal{F^{+}}$ be the set of non-negative integer solutions of the System of equations 1. Then, there exists a finite generating set over $\mathbb{N}$ of semiflows in $\mathcal{F^{+}}$.
The question of the existence of a finite generating set being solved for $\mathbb{N}$, it is necessarily solved for $\mathbb{Q^{+}}$ and $\mathbb{Q}$. This lemma is necessary not only to prove the decomposition theorem but also to claim the computability of the extremums described in Theorem 5.2.
Theorem 4.1
(Decomposition over $\mathbb{N}$) A semiflow is minimal if and only if it belongs to all generating sets over $\mathbb{N}$.
The set of minimal semiflows of $\mathcal{F^{+}}$ is a finite generating set over $\mathbb{N}$.
Let’s consider a semiflow $f\in\mathcal{F^{+}}\setminus\{0\}$ and its decomposition over any family of $k$ non-null semiflows $f_{i},1\leq i\leq k$. Then, there exist $a_{1},...,a_{k}\in\mathbb{N}$ such that $f=\sum_{i=1}^{i=k}a_{i}f_{i}$. Since $f\neq 0$ and all coefficients $a_{i}$ are in $\mathbb{N}$, there exists $j\leq k$ such that $0<f_{j}\leq a_{j}f_{j}\leq f$. If $f$ is minimal, then $a_{j}=1$ and $f_{j}=f$. Hence, if a semiflow is minimal, then it belongs to any generating set over $\mathbb{N}$. The reverse will become clear once the second statement of the theorem is proven.
Applying Gordan’s lemma, there exists a finite generating set, $\mathcal{G}$. Since any minimal semiflow is in $\mathcal{G}$, the subset of all minimal semiflows is included in $\mathcal{G}$ and therefore finite. Let =$\mathcal{E}=\{e_{1},...e_{n}\}$ be this subset and prove by construction that $\mathcal{E}$ is a generating set.
For any semiflow $f\in\mathcal{F^{+}}$, we build the following sequence leading to the decomposition of $f$:
i) $r_{0}=f$;
ii) $r_{i}=r_{i-1}-k_{i}e_{i}$ such that $r_{i}\in\mathcal{F^{+}}$ and $r_{i-1}-(k_{i}+1)e_{i}\notin\mathcal{F^{+}}.$
By construction of the non-negative integers $k_{i}$, we have $r_{n}\in\mathcal{F^{+}}$ and there doesn’t exist $e_{i}\in\mathcal{E}$ such that $e_{i}\leq r_{n}$. This means that $r_{n}$ is either minimal or null. Since $\mathcal{E}$ includes all minimal semiflows, therefore $r_{n}=0$, and any semiflow can be decomposed as a linear combinations of minimal semiflows; in other words, $\mathcal{E}$ is a finite generating set [^2] If $\mathcal{E}$ were to be infinite, the construction could still be used, since the monotonically decreasing sequence $r_{i}$ is bounded by 0 and $\mathbb{N}$ is nowhere dense, so we would have $\lim\limits_{n\to\infty}(f-\sum_{j=1}^{j=n}k_{j}e_{j})=0$, with the same definition of the coefficients $k_{j}$ as in ii).. It is now clear that if a semiflow $f$ belongs to any generating set, then it belongs in particular to $\mathcal{E}$; therefore, $f$ is a minimal semiflow. $\square$
Let’s point out that since $\mathcal{E}$ is not necessarily a basis, the decomposition is not unique in general and depends on the order in which the minimal semiflows of $\mathcal{E}$ are considered to perform the decomposition. However, a minimal semiflow does not necessarily belong to a generating set over $\mathbb{Q^{+}}$ or $\mathbb{Q}$.
4.2.2 Decomposition over semiflows of minimal support
These two theorems can already be found in [Mem23].
Theorem 4.2
(Minimal support) If $I$ is a minimal support, then
i) there exists a unique minimal semiflow $f$ such that $I=\left\|f\right\|$ and, for all $g\in\mathcal{F^{+}}$ such that $\left\|g\right\|=I$, there exists $k\in\mathbb{N}$ such that $g=kf$, and
ii) any non-null semiflow $g$ such that $\left\|g\right\|=I$ constitutes a generating set over $\mathbb{Q^{+}}$ or $\mathbb{Q}$ for $\mathcal{F}_{I}^{+}=\left\{g\in\mathcal{F^{+}}\>|\ \left\|g\right\|=I\right\}$.
In other words, $\{f\}$ is a unique generating set over $\mathbb{N}$ for $\mathcal{F}_{I}^{+}=\{g\in\mathcal{F^{+}}\;|\left\|g\right\|=I\}$. Indeed, this uniqueness property is lost in $\mathbb{Q^{+}}$ or in $\mathbb{Q}$, since any element of $\mathcal{F}_{I}^{+}$ is a generating set of $\mathcal{F}_{I}^{+}$ over $\mathbb{Q^{+}}$ or $\mathbb{Q}$.
Theorem 4.3
(Decomposition over $\mathbb{Q}^{+}$) Any support $I$ of semiflows is covered by the finite subset $\{I_{1},I_{2},\dots,I_{N}\}$ of minimal supports of semiflows included in $I$: $I=\bigcup_{i=1}^{i=N}I_{i}$.
Moreover, for all $f\in\mathcal{F^{+}}$ such that $\left\|f\right\|\subseteq I$, one has $f=\sum_{i=1}^{i=N}\alpha_{i}g_{i}$, where, for all $i\in\{1,2,...N\},\ \alpha_{i}\in\mathbb{Q^{+}}$ and the semiflows $g_{i}$ are such that $\left\|g_{i}\right\|=I_{i}$.
A sketch of the proof of Theorem 4.3 using Property 4 can be found in [MR79], and a complete proof, in [Mem78].
5 Home spaces and home states
The notion of home space was first defined in [Mem83] for Petri Nets relatively to a single initial marking. Here, we effortlessly extend its definition relatively to a nonempty subset of markings (or states if we were to consider Transition Systems).
Home spaces are extremely useful to analyze liveness or resilience (see [FH24]). Any behavioral property requiring to eventually become satisfied after executing a known sequence of transitions can be supported by a home space (a property satisfied for any reachable marking would be an invariant).
5.1 Definitions and basic properties
Given a Petri Net $PN$, its associated set $Q$ of all potential markings and a subset $Init$ of $Q$, we say that a set HS is an Init-home space if and only if, for any progression (i.e. sequence of transitions) from any element of $Init$, there exists a way of prolonging this progression and reach an element of HS. In other words:
Definition 5 (Home space)
[…]
This definition is general and can be applied to any Transition System. In [JL22], we can find, for Petri Nets, an equivalent definition: HS is an Init-home space if and only if $RS(M,Init)\subseteq RS^{-1}(M,HS\cap Q)$.
Definition 6 (Home state)
Given a nonempty subset Init of $Q$, a marking $s$ is an Init-home state if and only if $\{s\}$ is an Init-home space.
If $s$ is an Init-home state, then it is straightforwardly an $\{s\}$-home state, and we simply say that $s$ is a home state when there is no ambiguity. This is the usual notation that can be found in [Bra82], p.59, or in [GV03], p. 63, for Petri Nets. It can be found in many other papers such as [HDK14].
In many systems, the initial marking $q_{0}$ represents an idle state from which the various capabilities of the system can be executed. In this case, it is important for $q_{0}$ to be a home state. This property is usually guaranteed by a reset function that can be modeled in a simplistic way by adding a transition $r$ such that $\forall q\in RS(M,q_{0}),\ q_{0}\in r(q)$ (which means that $r$ is executable from any reachable marking and that its execution reaches $q_{0}$). However, by requiring to add too much complexity to $RG$ (one edge per node), this function is most of the time abstracted away when building $RG$ up.
It is not always easy to prove that a given set is an Init-home space. This question is addressed in [JL22] and is proven decidable for home state for Petri Nets but is still open in a more complex conceptual model. Furthermore, a corpus of decidable properties can be found in [FEJ89, FH24], or [JL22]. It may be worth mentioning the straightforward following properties, given two subsets $A$ and $B$ of markings.
Property 5
Any set containing an A-home space is also an A-home space. If $HS$ is an A-home space, it is a B-home space for any nonempty subset $B$ of $A$. If $HS_{1}$ is an $A_{1}$-home space and $HS_{2}$ is an $A_{2}$-home space, then $HS_{1}\cup HS_{2}$ is an ($A_{1}\cup A_{2}$)-home space.
However, the intersection of two home spaces is not necessarily a home space. Figure 1 represents the reachability graph of a transition system with eight markings. $HS_{1},HS_{2}$ and $HS_{3}$, as defined Figure 1, are three $\{q_{0}\}$-home spaces. While $HS_{1}\cap HS_{3}=\{q_{1},q_{3}\}$ is a $\{q_{0}\}$-home space, $HS_{1}\cap HS_{2}=\{q_{1}\}$ is not a $\{q_{0}\}$-home space (even if it is a $\{q_{1}\}$-home state).
Figure 1: With $Q=\{q_{0},q_{1},q_{2},q_{3},q_{4},q_{5},q_{6},q_{7}\},\ HS_{1}=\{q_{1},q_{3},q_{4}\},\ HS_{2}=\{q_{1},q_{5}\}$ and $\ HS_{3}=\{q_{1},q_{3},q_{5}\}$ are three $\{q_{0}\}$-home space. $HS_{4}=\{q_{1},q_{4},q_{7}\}$ is a $\{q_{6}\}$-home space as well as a $\{q_{0}\}$-home space.
Given a Petri Net $PN$ and a subset of markings $Init$, a sink is a marking with no successor in the associated reachability graph $RG(PN,Init)$. More generally, a subset $S$ of markings is a sink in $RG(PN,Init)$ if and only if $RS(M,S)=S$. Similarly, we say that strongly connected component $S$ of $RG(PN,Init)$ is strongly connected component sink if and only if $\nexists\ y\in RS(M,q_{0})\setminus S$ such that $\exists\ x\in S$ and $x\rightarrow y$. As any directed graph, $RG(M,Init)$ can have its vertices (markings) partitioned into strongly connected components and some of them can be sink at the same time. The following property is new and can be easily deduced from the definition of sink, strongly connected component, and home space.
Property 6
If there exists a unique strongly connected component sink $S$ in $RG(M,Init)$ then $S$ is a home space. Moreover, a marking is a home state if and only if it belongs to $S$. More generally, any home space has at least one element in each strongly connected component sink of the reachability graph [^3] It is easy to prove that this property holds even as the reachability graph can be infinite, considering that the definitions of sources, sinks, or strongly connected components are the same as in the case where the directed reachability graph is finite..
For the following property, we consider a Petri Net $PN$ paired with a single initial marking $q_{0}$.
Property 7
The three following statements are equivalent:
  • (i)
    the initial marking is a home state;
  • (ii)
    every reachable marking is a home state;
  • (iii)
    the reachability graph is strongly connected.
If $q_{0}$ is the initial marking, then, for all $x,y\in RS(PN,q_{0})$, there exists a path from $q_{0}$ to $x$ and a path from $q_{0}$ to $y$, and since $q_{0}$ is a home state, there also exists a path from $x$ to $q_{0}$ and from $y$ to $q_{0}$ in the reachability graph. Hence, $q_{0},x$ and $y$ belong to the same strongly connected component. We easily conclude that the reachability graph is strongly connected. The other elements of the property become obvious. $\square$
The strong connectivity of a given reachability graph means that some transitions are live. This remark suggests exploring this further in the following subsection.
5.2 Home spaces, semiflows, and liveness
Semiflows are intimately associated with home spaces and invariants and can greatly simplify the proof of fundamental properties of Petri Nets (even including parameters as in [BEI+20]) such as safeness, boundedness, or more complex behavioral properties such as liveness. Let us provide three properties supporting this idea.
Let Dom(t) denote the subset of markings from which the transition $t$ is executable, and Im(t), the subset of markings that can be reached by the execution of $t$.
Property 8
A transition $t$ is live if and only if Dom(t) is a home space.
Moreover, if Dom(t) is a home space, then Im(t) is also a home space.
This can be directly deduced from the usual definition of liveness and Definition 5 of home spaces. $\square$
We consider $\langle PN,q_{0}\rangle$, a Petri Net $PN$ with its initial marking $q_{0}$, its associated reachability set $RS$, its labeled reachability graph $LRG$, a home space $HS$ and $H=HS\cap RS$ such that $H$ induces (see, for instance, [Die17] for the notion of induced subgraph) a strongly connected subgraph of $LRG$.
Lemma 2
If a home space $H$ induces a strongly connected subgraph of LRG, then a transition $t$ is live if and only if there exist $h_{t}\in H$ and $\sigma\in T^{*}$ such that $h_{t}\xrightarrow{\overset{\sigma t}{}}$.
If $HS$ is a home space, then $H$ is also a home space, and for all $q\in RS$, there exist $s_{1}\in T^{*}$ and $h\in H$ such that $q\overset{s_{1}\ }{\rightarrow}h$.
The subgraph induced by $H$ being strongly connected, there exists a path from $h$ to $h_{t}$; in other words, there exists $s_{2}\in T^{*}$ such that $h\overset{s_{2}\ }{\rightarrow}h_{t}$. We can construct a sequence $s=s_{1}s_{2}\sigma$ such that for all $q\in RS,q\overset{st}{\rightarrow}$. Hence $t$ is live in $RS(M,q_{0})$. The reverse is obvious. $\square$
From this lemma, we can easily deduce the following property regarding home states:
Property 9
Let $PN$ be a Petri Net and $q$ be a home state. Then, any transition that is enabled at $q$ is live, and, more generally, a transition is live if and only if it appears as a label in $LRG(PN,q)$.
This can easily be proven directly from the definition of liveness and Lemma 2 about home states $\square$
We can then deduce from this property that liveness is decidable for Petri Nets equipped with a home state. More precisely, we have:
Theorem 5.1
Let PN be a Petri Net with a home state $q$, and $LCT(PN,q)$, the coverability tree of PN. A transition is live if and only if it appears as a label in $LCT(PN,q)$.
This can be proven directly from the fact that a transition appears as a label of an edge of $LRG(PN,q)$ if and only if it appears as a label of an edge of $LCT(PN,q)$, and by considering Property 9. $\square$
Corollary 2
For any Petri Net with a home state $q$, liveness is decidable.
This is a direct consequence of Theorem 5.1 combined with Karp and Miller’s theorem [KM69], stating that the coverability tree is finite and considering $\square$
Given an initial state $q_{0}$, each semiflow can be associated with an invariant that, in turn, can be associated with a home space. In other words, if $f\in\mathcal{F}$, then $HS(f,q_{0})=\{q\in Q\ |\ f^{\top}q=f^{\top}q_{0}\}$ is a $\{q_{0}\}$-home space, since $RS(M,q_{0})\subseteq HS(f,q_{0})$.
Property 10
If $f\in\mathcal{F}$, then, for all $\alpha\in\mathbb{Q}\setminus\{0\},HS(\alpha f,q_{0})=HS(f,q_{0})$. Also, for all $f$ and $g\in\mathcal{F}$ and for all $\alpha$ and $\beta\in\mathbb{Q},HS(f,q_{0})\cap HS(g,q_{0})\subseteq HS(\alpha f+\beta g,q_{0})$. Moreover, $HS(f,q_{0})\cap HS(g,q_{0})$ is a $\{q_{0}\}$-home space.
Note that $HS(f,q_{0})\ \cap\ HS(g,q_{0})$ is straightforwardly a $\{q_{0}\}$-home space, since they both contain $RS(M,q_{0})$ [^4] Let us recall that, in general, the intersection of home spaces is not a home space (see Figure 1).. If $q\in HS(f,q_{0})\cap HS(g,q_{0})$, then $\alpha(f^{\top}q)=\alpha(f^{\top}q_{0})$ and $\beta(g^{\top}q)=\beta(g^{\top}q_{0})$, so $(\alpha f+\beta g)^{\top}q=(\alpha f+\beta g)^{\top}q_{0}$, and, therefore, $q\in HS(\alpha f+\beta g,q_{0})$square
These results provide us with a few steps to analyze and prove that a subset of transitions are live. From a set of invariants, we can define a first home space $HS$ that concisely describes how tokens are distributed over places. From this token distribution, we can analyze what transition are enabled in order to prove that a specific given marking $q$ ($q_{0}$ being the usual case) is always reachable from any element of $HS$. When this is possible, it can easily be deduced that $q$ is a home state. Then, by using Theorem 5.1, it can be proven which transitions are live and, ultimately, whether the Petri Net is live or not. This will be illustrated later with examples in Section 6.
5.3 Three extremums drawn from the notion of semiflow
The knowledge of any finite generating set allows a practical computation of three extremums directly inspired from Property 5 (Section 5.2) and Corollary 1 (Section 3). First, starting from an initial marking $q_{0}$, we can define $\mathcal{HS}$, the smallest subset of $Q$ stable for $\cap$ and generated by $\{H\subseteq Q\ |\ \exists f\in\mathcal{F}^{+},\ H=HS(f,q_{0})\}$.
Let us note that the reachability set is included in any element of $\mathcal{HS}$; therefore, there exists a unique nonempty least element of $\mathcal{HS}$.
Definition 7 (Extremums)
Given an initial marking $q_{0}$ and the set of semiflows $\mathcal{F^{+}}$, the three following extremums can be defined:
    $\iota=min(\mathcal{HS})$ is the least element of $\mathcal{HS}$;
    $\mu(p,q_{0})=\min_{\{f\in\mathcal{F}^{+}\ |\ f(p)\neq 0\}}\frac{f^{\top}q_{0}}{f(p)}$ is the lowest bound that can be built directly from a semiflow the support of which contains the given place $p$ in $P$;
    $\rho=max\{S\subseteq P\ |\ \exists f\in\mathcal{F}^{+},S=\left\|f\right\|\}$ is the largest support of any semiflow in $\mathcal{F^{+}}$.
Theorem 5.2 expresses the fact that these extremums are computable as soon as any generating set is available:
Theorem 5.2
Let =$\mathcal{E}=\{e_{1},...e_{N}\}$ be any finite generating set of $\mathcal{F}^{+}$, and $q_{0}\in Q$, an initial marking.
  • If $\mathcal{E}$ is over $\mathbb{S}$, then we have: $\iota=\bigcap_{f\in\mathcal{F^{+}}}HS(f,q_{0})=\bigcap_{e_{i}\in\mathcal{E}}HS(e_{i},q_{0})$;
  • If $\mathcal{E}$ is over $\mathbb{Q}^{+}$ or $\mathbb{N}$, then, for any place $p$ belonging to at least one support of a semiflow of $\mathcal{F}^{+}$, for all $q\in RS(PN,q_{0})$, we have :
    $q(p)\leq\mu(p,q_{0})=\min_{\{f\in\mathcal{F}^{+}\ |\ f(p)\neq 0\}}\frac{f^{\top}q_{0}}{f(p)}=\min_{\{e_{i}\in\mathcal{E}\ |\ e_{i}(p)\neq 0\}}\frac{{e_{i}}^{\top}q_{0}}{e_{i}(p)};$
  • If $\mathcal{E}$ is over $\mathbb{S}$, then we have: $\rho=\left\|\sum_{f\in\mathcal{F}^{+}}f\right\|=\bigcup_{f\in\mathcal{F}^{+}}\left\|f\right\|=\bigcup_{e_{i}\in\mathcal{E}}\left\|e_{i}\right\|$.
  • For the first item, let’s consider $f\in\mathcal{F^{+}}$ with $f=\sum_{i=1}^{i=N}\alpha_{i}e_{i}$ and $q\in\bigcap_{e_{i}\in\mathcal{E}}HS(e_{i},q_{0})$. Then, $\alpha_{i}(e_{i}^{\top}q)=\alpha_{i}(e_{i}^{\top}q_{0})$, for all $i\in\{1,...N\}$, and, hence $\sum_{i=1}^{i=N}\alpha_{i}(e_{i}^{\top}q)=\sum_{i=1}^{i=N}\alpha_{i}(e_{i}^{\top}q_{0})$. Then, for all $f\in\mathcal{F^{+}}$, $f^{\top}q=f^{\top}q_{0}$, and $q\in HS(f,q_{0})$.
    Therefore, since $\mathcal{E}\subset\mathcal{F^{+}}$ directly implies $(\bigcap_{f\in\mathcal{F^{+}}}HS(f,q_{0}))\subseteq\bigcap_{e_{i}\in\mathcal{E}}HS(e_{i},q_{0})$, we have: $\bigcap_{e_{i}\in\mathcal{E}}HS(e_{i},q_{0})=\bigcap_{f\in\mathcal{F^{+}}}HS(f,q_{0})=\iota$.
  • For the second item of the theorem, let’s consider a marking $q_{0}$, a place $p$, and a semiflow $f$ of $\mathcal{F}^{+}$ such that $f(p)>0$ and $f=\sum_{i=1}^{i=N}\alpha_{i}e_{i}$, where $\alpha_{i}\geq 0$, for all $i\in\{1,...N\}$.
    Let’s define $\mu_{\mathcal{E}}$ such that $\mu_{\mathcal{E}}=\min_{\{e_{i}\in\mathcal{E}\ |\ e_{i}(p)\neq 0\}}\frac{{e_{i}}^{\top}q_{0}}{e_{i}(p)}$. Then, there exists $j$ such that $1\leq j\leq N$ and $\mu_{\mathcal{E}}=\frac{{e_{j}}^{\top}q_{0}}{e_{j}(p)}$.
    Therefore, for all $i\leq N$ such that $e_{i}(p)\neq 0$, there exists $\delta_{i}\in\mathbb{Q}^{+}$ such that: $\frac{{e_{j}}^{\top}q_{0}}{e_{j}(p)}=\frac{{e_{i}}^{\top}q_{0}-\delta_{i}}{e_{i}(p)}$. It can then be deduced, for all such $i$:
    $\mu_{\mathcal{E}}=\frac{\alpha_{j}{e_{j}}^{\top}q_{0}}{\alpha_{j}e_{j}(p)}=\frac{\alpha_{i}({e_{i}}^{\top}q_{0}-\delta_{i})}{\alpha_{i}e_{i}(p)},$
    and, therefore:
    $\begin{array}[]{ll}\mu_{\mathcal{E}}&=\frac{\sum_{\{i\ |\ e_{i}(p)>0\}}\alpha_{i}({e_{i}}^{\top}q_{0}-\delta_{i})}{\sum_{\{i\ |\ e_{i}(p)>0\}}\alpha_{i}e_{i}(p)}\\ &=\frac{\sum_{\{i\ |\ e_{i}(p)>0\}}\alpha_{i}({e_{i}}^{\top}q_{0}-\delta_{i})+\sum_{\{i\ |e_{i}(p)=0\}}(\alpha_{i}{e_{i}}^{\top}q_{0}-\alpha_{i}{e_{i}}^{\top}q_{0})}{\sum_{\{i\ |\ e_{i}(p)>0\}}\alpha_{i}e_{i}(p)}\\ &=\frac{\sum_{\{i|e_{i}(p)>0\}}\alpha_{i}{e_{i}}^{\top}q_{0}+\sum_{\{i|e_{i}(p)=0\}}\alpha_{i}{e_{i}}^{\top}q_{0}-\sum_{\{i|e_{i}(p)>0\}}\alpha_{i}\delta_{i}-\sum_{\{i\ |\ e_{i}(p)=0\}}\alpha_{i}{e_{i}}^{\top}q_{0}}{\sum_{\{i\ |\ e_{i}(p)>0\}}\alpha_{i}e_{i}(p)+\sum_{\{i\ |\ e_{i}(p)=0\}}\alpha_{i}e_{i}(p)},\end{array}$
    since $\sum_{\{i\ |\ e_{i}(p)=0\}}\alpha_{i}e_{i}(p)=0$. Then, since $\delta_{i}\geq 0$ and $\alpha_{i}\geq 0$ for all $i$ such that $1\leq i\leq N$,
    $\begin{array}[]{ll}\mu_{\mathcal{E}}&=\frac{\sum_{i=1}^{i=N}\alpha_{i}{e_{i}}^{\top}q_{0}-\sum_{\{i|e_{i}(p)>0\}}\alpha_{i}\delta_{i}-\sum_{\{i\ |\ e_{i}(p)=0\}}\alpha_{i}{e_{i}}^{\top}q_{0}}{\sum_{i=1}^{i=N}\alpha_{i}e_{i}(p)}\\ &=\frac{f^{\top}q_{0}-\sum_{\{i|e_{i}(p)>0\}}\alpha_{i}\delta_{i}-\sum_{\{i\ |\ e_{i}(p)=0\}}\alpha_{i}{e_{i}}^{\top}q_{0}}{f(p)}\\ &\leq\frac{f^{\top}q_{0}}{f(p)}.\end{array}$
    This being verified for any semiflow of $\mathcal{F}^{+}$, we have $\mu(p,q_{0})=\mu_{\mathcal{E}}$.
  • For the third item of the theorem, let’s consider $\mathcal{E}$, a generating set over $\mathbb{S}$. Then, any semiflow $f$ in $\mathcal{F}^{+}$ can be decomposed as follows:
    $f=\sum_{\alpha_{i}>0}\alpha_{i}e_{i}+\sum_{\alpha_{i}<0}\alpha_{i}e_{i},$
    where $\alpha_{i}\in\mathbb{S}$. Since $f\in\mathcal{F}^{+}$, this means that at least one coefficient $\alpha_{i}$ is strictly positive, and $\sum_{\alpha_{i}<0}|\alpha_{i}|e_{i}+f=\sum_{\alpha_{i}>0}\alpha_{i}e_{i}\neq 0$.
    Therefore, applying Property 4:
    $\left\|f\right\|\subseteq\left\|\sum_{\alpha_{i}<0}|\alpha_{i}|e_{i}+f\right\|=\left\|\sum_{\alpha_{i}>0}\alpha_{i}e_{i}\right\|=\bigcup_{\alpha_{i}>0}\left\|\alpha_{i}e_{i}\right\|\subseteq\bigcup_{e_{i}\in\mathcal{E}}\left\|e_{i}\right\|.$
    Hence, $\rho=\left\|\sum_{f\in\mathcal{F}^{+}}f\right\|=\bigcup_{e_{i}\in\mathcal{E}}\left\|e_{i}\right\|$. $\square$
This theorem means that these three extremums, $\iota$, $\mu$ and $\rho$, can be computed with the help of one finite generating set. The third part of this theorem means that, if =$\mathcal{E}=\{e_{1},...e_{N}\}$ is any generating set of a given Petri Net PN, then $\bigcup_{e_{i}\in\mathcal{E}}\left\|e_{i}\right\|$ is also the unique largest support of PN.
6 Reasoning with invariants, semiflows, and home spaces
Invariants, semiflows, and home spaces can be used to prove a rich array of behavioral properties of Petri Nets, even within different settings, in particular when using parameters.
[…]
6.1 A tiny example
The Petri Net $TN(i)=\langle\{A,B\},\{t_{1},t_{2}\},Pre,Post\rangle$ in Figure 2 is defined by:
$Pre(\cdot,t_{1})^{\top}=(i,0);Pre(\cdot,t_{2})^{\top}=(1,1)$;
$Post(\cdot,t_{1})^{\top}=(0,1);Post(\cdot,t_{2})^{\top}=(i+1,0).$
We consider the initial marking $q_{0}$ such that $q_{0}(A)=n$ and $q_{0}(B)=x$, where $n$ and $x\in\mathbb{N}$.
Figure 2: Semiflows must verify the equation $i\times a=b$, for which $g^{\top}=(1,i)$ is an obvious solution. $TN(i)$ is live if and only if $g^{\top}q_{0}>i$ and is not a multiple of $i$, regardless of the initial marking of $B$. For $i=1$, $TN(1)$ has no live transition, regardless of the initial marking.
A first version of this example can be found for $i=2$ in [Bra82] or in [Mem83], without proof. Here, the Petri Net is enriched by introducing a parameter $i$ such that $i>1$.
We have the following minimal semiflow of minimal support: $g^{\top}=(1,i)$, and we can prove that $\left\langle TN(i),q_{0}\right\rangle$ is not live if and only if $g^{\top}q_{0}\leq i$ or $g^{\top}q_{0}=n\times i$, independently of the value $x$ of $q_{0}(B)$. In other words, $TN(i)$ recognizes whether a given number $n$ is a multiple of $i$.
First, if $g^{\top}q_{0}<i$, then the enabling threshold of $t_{1}$ can never be reached (Property 3) and neither $t_{1}$ nor $t_{2}$ can be executed (since $q_{0}(B)$ is necessarily null to satisfy the inequality). Second, if $g^{\top}q_{0}\geq i$, then we consider the Euclidean division of $g^{\top}q_{0}$ by $i$, giving $g^{\top}q_{0}=n\times i+r$, where $r<i$. Then, since $g$ is a semiflow, $g^{\top}q=q(A)+iq(B)\equiv r\mod i$, and, therefore, $q(A)\equiv r\mod i$, for all $q\in RS(TN(i),q_{0})$. If $r=0$, then we have $q(A)=n\times i-i\times q(B)$, and $t_{1}$ can be executed $n-q(B)$ times to reach a marking with zero token in $A$.
If $r\neq 0$ and $g^{\top}q_{0}>i$, then $HS=\{q\in RS(TN(i),q_{0})\ |\ q(A)\neq 0\ \land\ q(B)\neq 0\}$ is a home space; therefore, it is always possible to execute $t_{2}$. It is easy to conclude that the Petri Net $TN(i)$ is live if and only if $g^{\top}q_{0}>i$ and is not a multiple of $i$, regardless of the initial marking of $B$. $\square$
What is remarkable about the analysis of this tiny example is that it was not necessary to develop a symbolic reachability graph in order to decide whether or not the Petri Net is live or bounded. We could analyze the Petri Net even partially ignoring the initial marking (i.e., considering $q_{0}(A)$ as an additional parameter and without even considering the value taken by $q_{0}(B)$).
6.1.1 Euclidean division
From the properties of $TN(i)$, it is natural to progress by one more step and propose to design a Petri Net with the ability not only to recognize whether a natural number $n$ is a multiple of a given natural number $i$, but more generally to recognize the remainder of the Euclidean division of $n$ such that $n>0$ by $i$ such that $i>1$. To this effect, we first consider the Colored Petri Net $TNCED(i)$ of Figure 3, and the parameter $i\geq 2$. Second, for easing the reasoning, we unfold $TNCED(i)$ into the classic Petri Net $TNED(i)$, where each place $A_{j}$ represents the color $j$ of $TNCED(i)$ (see Figures 3 and 4).
[…]
$Pre(A_{j},t_{j,1})=i,\ Pre(B,t_{j,1})=\ Pre(A_{j},t_{j,2})=1$
$Post(A_{j},t_{j,2})=i+1,Post(B,t_{j,1})=1.$
where $j\in\left[0,i-1\right]$. The initial marking is such that $q_{0}(A_{j})=n+j$, where $j\in\left[0,i-1\right]$, and $q_{0}(B)=x$, where $n>0$ and $x$ are natural numbers.
Figure 3: $TNCED(i)$, is a colored Petri Net with a set $C$ of colors representing any value of the parameter $i$ between 0 and $i-1$, and $\tau$ is an undefined token; $\ C=([0,...i-1]\cap\mathbb{N})\cup\{\tau\}$. This time, we have a system of $i$ equations: $i\times a_{j}=b$ with $j\in\left[0,i-1\right]$, for which $g$ such that $g(A_{j})=1$ for $j\in\left[0,i-1\right]$ and $g(B)=i$ is the minimal semiflow of minimal support in $\mathbb{N}$. This parameterized Colored Petri Net allows knowing the remainder of the Euclidean division of a natural number $n$ by $i$.
Figure 4: $TNED(i)$ is the unfolded version of $TNCED$. To compute semiflows, we have a system of $i$ equations: $i\times a_{j}=b$ with $j\in\left[0,i-1\right]$, for which $g$ such that $g(A_{j})=1$ for $j\in\left[0,i-1\right]$ and $g(B)=i$ is the minimal semiflow of minimal support in $\mathbb{N}$. This parameterized Petri Net allows knowing the remainder of the Euclidean division of a natural number $n$ by $i$.
We set $g_{i}^{T}=(1,\cdots 1,i)$, such that $g(A_{j})=1$ for $j\in\left[0,i-1\right]$, and $g(B)=i$ is the minimal semiflow of minimal support in $\mathbb{N}$. We have a first invariant $I$, for all $q\in RS(TNED(i),q_{0})$:
$g^{\top}q_{0}=g^{\top}q=\sum_{j=0}^{j=i-1}q_{0}(A_{j})+iq_{0}(B)=i\times(x+n+\frac{i-1}{2}).$
Then, we need to notice that any place $A_{j}$ is connected to only two transitions, $t_{j,1}$ and $t_{j,2}$, such that:
$Post(A_{j},t_{j,1})-Pre(A_{j},t_{j,1})=-i$,
$Post(A_{j},t_{j,2})-Pre(A_{j},t_{j,2})=i$.
Hence, for all $q\in RS(TNED(i),q_{0})$ and $j\in\left[0,i-1\right]$, $q(A_{j})$ can only vary by ±$\pm i$. We then deduce a family of invariants $I(j)$ for $j\in\left[0,i-1\right]$:
$I(j):\forall q\in RS(TNED(i),q_{0}),\ q(A_{j})\equiv q_{0}(A_{j})\mod i.$
Let’s perform the Euclidean division of $q_{0}(A_{j})$ by $i$. We have: $q_{0}(A_{j})=n+j=a_{j}\times i+\alpha_{j}$, where $\alpha_{j}<i$ for all $j\in\left[0,i-1\right]$. A new family of invariants $I^{\prime}(j)$ can be directly deduced from each $I(j)$:
$I^{\prime}(j):\forall q\in RS(TNED(i),q_{0}),\ q(A_{j})\geq\alpha_{j}.$
Furthermore, it must be pointed out that $\{\alpha_{0},\cdots\alpha_{i-1}\}$ is a permutation of $\{0,\cdots i-1\}$. Indeed, if there exist $j<i$ and $j^{\prime}<i$ such that $\alpha_{j}=\alpha_{j^{\prime}}$, then $n+j-a_{j}\times i=n+j^{\prime}-a_{j^{\prime}}\times i$, and $|j-j^{\prime}|=|a_{j}-a_{j^{\prime}}|\times i$. Since $|j-j^{\prime}|<i$, we have $a_{j}=a_{j^{\prime}}$ and $j=j^{\prime}$. Therefore,
  • (a)
    $\sum_{j=0}^{j=i-1}q_{(}A_{j})\geq\frac{i(i-1)}{2}$ (directly from the $I^{\prime}(j)$ family of invariants), and
  • (b)
    there is a unique $k\in\left[0,i-1\right]$ such that $\alpha_{k}=0$.
From (a) and $I$, we deduce, for all $q\in RS(TNED(i),q_{0}),\ q(B)\leq x+n$ (which is a better bound than the one that can be deduced from Proposition 2). Also, from $I(j)$, we can deduce, for ll $q\in RS(TNED(i),q_{0}),q(A_{j})=y_{j}\times i+\alpha_{j},$ where $y_{j}\in\mathbb{N}$.
From any reachable marking $q$, the sequence $\sigma_{q}=t_{0,1}^{y_{0}}\cdots t_{i-1,1}^{y_{i-1}}$ can be executed and reach the marking $q_{h}$ such that, for all $j\in\left[0,i-1\right],\ q_{h}(A_{j})=\alpha_{j}$ and $q_{h}(B)=x+n$.
We know $q_{h}$ is a home state, since $\sigma_{q}$ is defined for any reachable marking (Note that $q_{0}$ is not a home state). From Property 9, we deduce that, since any transition $t_{j,2}$ where $j\neq k$ is executable ($n>0$ hence, $q_{h}(B)>1$ and $q_{h}(A_{j})=\alpha_{j}>0$), then $t_{j,2}$ is live, and, therefore, the corresponding transitions $t_{j,1}$ are also live.
From (b), we have $q_{h}(A_{k})=0$ from, which we deduce that $t_{k,1}$ and $t_{k,2}$ are not live [^5] Actually, it suffices to notice that $A_{k}$ is an empty deadlock that remains empty.. Finally, we have $n+k=a_{k}\times i$, and the remainder of the Euclidean division of $n$ by $i$ is $i-k$.
$TNED(i)$ provides the ability to recognize this remainder by the remarkable fact that $(t_{k,1},t_{k,2})$ is the only couple of transitions not live $\square$
Most of the time, in real-life use cases, when a model accepts a set of home states, then the initial marking belongs to it. It is not the case in our example, and this suggests the following conjecture:
“If the initial marking $q_{0}$ of a given Petri Net $PN$ is not a home state and there exists a home state reachable from $q_{0}$, then there exists at least one non-live transition in $\langle PN,q_{0}\rangle$."
7 Conclusion
It has been recalled how semiflows create a link from the static structure (i.e., the bipartite graph) to the dynamic evolution (i.e., the variation of the number of tokens) of Petri Nets. They support constraints over all possible markings, which greatly help analysis of behavioral properties (even the discovery of some unspecified ones). More generally, analysis can be performed with incomplete information, particularly when markings are described with parameters as in our two examples. Petri Nets transitions can also be described with parameters (as in the Colored Petri Net of our second example) that are making the invariant calculus more complex but still tractable. Most of the time, especially with real-life system models, it will be possible to avoid a painstaking symbolic model checking or a parameterized and complex development of a reachability graph [DRvB01].
Semiflows infer a class of invariants that can be deduced by algebraic computation. Furthermore, the set of semiflows can be characterized with the notion of minimal generating set, and we hope that our three decomposition theorems reached their final version.
We introduced new results about home spaces; in particular, theorem 5.1 is new to the best of our knowledge (for instance, it does not appear in the recent survey on decidability issues for Petri Nets [EN24]). This theorem is interesting for at least two different reasons. First, from a theoretical point of view, since the existence of a home state does not mean that the Petri Net is bounded. Second, from a practical point of view, since systems often have a home state by design. It increases the importance one can grant to the construction of coverability trees, which is used mostly to determine which places are bounded (see important works by Finkel and al [FHK20] about accelerating this construction).
Theorem 5.2 is an indication that a generating set brings most of the information that semiflows in $\mathcal{F}^{+}$ can provide for analysis of behavioral properties.
At last, we presented most of these results in the framework of Petri Nets, we believe that most of these results apply to Transition Systems. This, indeed, constitutes a starting point for future work.